OpenStudy (anonymous):

compute the first and second derivative of the function f(x)=(5x^2 +x)^4

myininaya (myininaya):

f(x)=g(h(x)) => f'(x)=h'(x)g'(h(x)) => f''(x)=h''(x)g'(h(x))+h'(x)h'(x)g''(h(x))

OpenStudy (anonymous):

okay so is f'(x)=4(5x^2+x)^3(10x)+1

myininaya (myininaya):

almost the paranthesis should be around 10x+1

myininaya (myininaya):

not just the 10x

OpenStudy (anonymous):

okay so f'(x)=4(5x^2+x)^3 (10X+1)

myininaya (myininaya):

yes except that X should be a little x but you knew that

OpenStudy (anonymous):

yes, thank you. for f"(x) I got 12(5x^2+x)^2 (10x+1) (10), is this correct

myininaya (myininaya):

you will need product rule to find the derivative of f'

myininaya (myininaya):

we have h' times g'(h(x)) applying product rule: h'' times g'(h(x)) + h' times h' g''(h(x)) /\ | here I used chain rule to find the derivative of g'(h(x)) you remember derivative of inside times derivative of outside

OpenStudy (anonymous):

okay

myininaya (myininaya):

\[f'=[4(5x^2+x)^3](10x+1)\] \[f''=[4(5x^2+x)^3]'(10x+1)+[4(5x^2+x)^3](10x+1)'\]

myininaya (myininaya):

\[[4(5x^2+x)^3]'=4 \cdot 3 (5x^2+x)^{3-1} (5x^2+x)'=12(5x^2+x)^2(10x+1)\]

myininaya (myininaya):

\[(10x+1)'=(10x)'+(1)'=10+0\]

OpenStudy (anonymous):

okay, thank you very much

myininaya (myininaya):

no problem you understand to plug those last two lines of mine into the expression i wrote for f'' right?

OpenStudy (anonymous):

yes, but would for [4(5x^2+x)^3]' would it be 12(5x^2+x)^2 (10x+1) (10) or not

myininaya (myininaya):

\[[(5x^2+x)^n]'=n(5x^2+x)^{n-1}(5x^2+x)' \] so no

myininaya (myininaya):

\[[h(g(x))]'=h'(g(x)) \cdot g'(x)\]

myininaya (myininaya):

the outside function is h(x)=x^3 and inside function is g(x)=5x^2+x h'(x)=3x^2 g'(x)=10x+1 h'(g(x))=3(5x^2+x)^2 so \[[h(g(x))]'=[(5x^2+x)^3]'=3(5x^2+x)^2(10x+1)\]

OpenStudy (anonymous):

okay, thanks

myininaya (myininaya):

np have a good day

OpenStudy (anonymous):

thanks, you too